# Pulsating with discrete symmetry

### v2

@@ Abstract
We propose a class of diffusive lattice models where the internal states of particles are subject to a periodic drive. The discrete symmetry of states enforces an effective energy landscape which can counteract the drive to arrest the dynamics. Such a mechanism is reminiscent of the arrest reported in dense assemblies of pulsating deformable particles, in line with experiments in confluent biological tissues. We show that the competition between arrest and synchronisation promotes an instability which leads to spiral waves. Using analytical coarse-graining, we rationalize the emergence of such spirals in terms of a discrete gauge invariance at the hydrodynamic level.
@@
Active matter describes non-equilibrium systems where a constant energy injection at the particle level leads to collective dynamics far from equilibrium, as is the case for many chemical, biological, and robotic systems [1, 2, 3, 4]. In the last decades, most studies of active matter have focused on the role of self-propulsion, i.e. the ability of each particle to independently undergo directed motion. This paradigm has led to the theoretical understanding of several kinds of collective dynamics which have no counterpart in equilibrium systems [5, 6].
@@
Interestingly, the energy injection in complex units need not result only into a self-propulsion mechanism. An important example is the case of deformable particles [7]. Indeed, complex aggregates such as macromolecules or living cells can change their shape due to internal activity [8], which leads to the spontaneous propagation of contraction waves in dense tissues. Such a wave propagation plays a crucial role in morphogenesis [9, 10, 11, 12, 13, 14], uterine contraction [15, 16], and cardiac arrhytmogenesis [17, 18].
@@
Vertex models are popular to capture the behavior of dense active systems such as biological tissues [19]. They typically consider self-propulsion as the only active component [20, 21]. Yet, when dense tissues behave like solids, it is questionable whether self-propulsion should be the key ingredient. Other models have considered dense assemblies of active deformable particles [22, 23, 24, 25, 26, 27], where energy injection here occurs through the sustained oscillation of size at the single-particle level. With a Kuramoto-like synchronisation [28, 29] of particle sizes, contraction waves spontaneously emerge [26, 27], which are reminiscent of those reported in biological systems [13, 17].
@@
Contraction waves in deformable active particles stem from the competition between synchronisation and steric repulsion [27]. The former favors a global cycling of particle sizes, while the latter promotes an arrested state without any cycling. Combining both stabilizes contraction waves. Interestingly, there are alternative mechanisms for arrest and synchronisation beyond steric repulsion and Kuramoto-like interactions. It remains to explore whether these alternatives lead to a phenomenology similar to that of deformable active particles. Indeed, one may wonder whether the arrest-synchronisation competition actually is a generic scenario for pattern formation. If so, it encourages one to search for any hydrodynamic invariance and/or broken symmetry which might stand out as a hallmark of this competition.
Figure 1: (a) Particles freely diffuse without any volume exclusion. (b) Each particle is subject to an independent periodic driving which pulsates its internal state. Fully-connected onsite interactions locally synchronise the pulsation of particles. (c-d) Instabilities promote the coexistence of pulsating domains. Rotating defects appear at the meeting point of domain interfaces, thus forming spiral waves with a threefold symmetry. $(L,f,\varepsilon,D_{0},\rho_{0})=(512,1.25,2.5,10^{3},10)$.
@@
In this Letter, we formulate a class of pulsating lattice models (PLM) featuring spiral waves with rotating defects [Fig. 1]. We reveal that the discrete symmetry of available states is key to arrest the dynamics, which, in combination with synchronisation and drive, generically yields wave propagation. Through analytical coarse-graining, we map PLM into a specific form of the complex Ginzburg-Landau equation (CGLE) [30] which breaks the continuous gauge invariance, while maintaining a discrete gauge invariance. We demonstrate how such an invariance not only signals the emergence of an arrested phase, but also constrains the features of the waves, at both the hydrodynamic and the particle-based levels. Overall, the arrest-synchronisation competition in PLM provides a generic route for pattern formation, distinct from the well-studied instabilities in other models of reaction-diffusion systems (RDS) [31, 32, 33].
@@
Pulsating lattice models.—We consider $N$ particles in $V=L^{d}$ sites of a hypercubic $d$-dimensional lattice, with global number density $\rho_{0}=N/V$, without any excluded volume. Each site contains an arbitrarily large number of particles. Each particle has an internal state, labeled by a discrete index $a\in\{1,\ldots,q\}$, as a proxy to mimic the internal phase of deformable particles [22, 23, 24, 25, 26, 27]. The crucial difference is that such states now feature a discrete symmetry. In what follows, we focus on the case $q=3$, which is the minimal number of states to accommodate a current, and $d=2$. The system configuration $\{n_{{\bf j},a}\}$ is then given by the number of particles for each state $a$ and site ${\bf j}$.
@@
At every time step $dt/N$, a particle with state $a$ can either jump to a neighbouring site with probability $D_{0}\,dt$, or switch to state $b$ with probability $W_{ba}\,dt$ [Figs. 1(a,b)]:
@@
The first term in the exponent of Eq. (1) is defined by $f_{ab}=\pm f$ when $a-b=\pm 1\mod 3$, and $f_{ab}=0$ otherwise. This drive mimics the pulsation of deformable active particles [22, 23, 24, 25, 26, 27]. Inspired by some active lattice dynamics [34, 35, 36, 37, 38, 39, 40], the second term accounts for a synchronising Potts-like [41] interaction with strength $\varepsilon$, where $\rho_{\bf j}=\sum_{a}n_{{\bf j},a}$ is the local density at site $\bf j$. This term favors transition towards the state with the highest number of particles locally.
@@
While interactions between particles are fully connected onsite, different sites exchange information only via diffusion. At small $D_{0}$, many transitions of internal states occur in between two rare jumps of sites. At large $D_{0}$, jumps are so frequent that now all particles are effectively interacting in between two state transitions. In both cases, patterns cannot emerge, and the system can be described solely in terms of the onsite dynamics [42, 43].
Synchronisation and arrest in the onsite case.—In the absence of diffusion ($D_{0}=0$), the evolution of the onsite occupation numbers $n_{a}$ can been studied in terms of the complex parameter
$A(t)=\frac{1}{\rho_{0}}\sum_{a=1}^{3}e^{\frac{2\pi i}{3}a}n_{a}(t).$ (2)
@@
There are three stable phases [42, 43]: (i) a disordered phase at small $\varepsilon$, where particles are uniformly distributed in the three states, with symmetric fixed point $n_{a}=\rho_{0}/3$, and $|A|$ vanishes at large $\rho_{0}$; (ii) a cycling phase at intermediate $\varepsilon$, where particles collectively undergo periodic transitions between states; (iii) an arrested phase at large $\varepsilon$, with three asymmetric fixed points $n_{a}\gg\rho_{0}/3$ invariant under the cyclic permutations $A\to A\,e^{\frac{2\pi i}{3}k}$ ($k$ integer). Transitions between states can here be regarded as unimolecular reactions, in contrast with some models of multimolecular RDS [44, 45]. Besides, the total number of particles here is conserved. These features lead to stabilising disorder, and ensure that the arrest is not an absorbing phase.
In the cycling phase, the discrete nature of the internal states enforces that the global phase $\arg(A)$ is subject to an effective energy landscape. Indeed, the local extrema of the Potts-like interaction, at $\arg(A)=k\pi/3$, constrain the dynamics of the phase. While Kuramoto oscillators (with continuous symmetry) synchronously cycle without any cost, the collective transition of Potts oscillators (with discrete symmetry) between two nearby synchronised configurations implies crossing a desynchronised configuration, which is not energetically favored. Consequently, as $\varepsilon$ increases, Potts oscillators spend more time in a given state before cycling to the next one. Above a critical $\varepsilon$, this energetic effect completely counteracts the drive, breaking down the periodicity of oscillations, and eventually stabilizing the arrested phase [46]. A similar phase trapping has been reported in clock models of aligning self-propelled particles [40].
@@
In short, arrest emerges in PLM solely due to the discrete symmetry of internal states. Therefore, PLM entail a competition between arrest and synchronisation despite the absence of any volume exclusion, in contrast with models of continuously deformable particles [26, 27]. This competition opens the door to the emergence of dynamical patterns in spatially extended systems.
@@
Spiral waves in the spatially extended case.—In the presence of diffusion ($D_{0}>0$), the displacement of particles follows a free dynamics, independently of any interaction, so that the density profile is always homogeneous. Yet, the spatial distribution of the particle states may not remain homogeneous. Indeed, even when particles are synchronised on site, they might not be synchronised between sites, so that the system can potentially accommodate spatial instabilities.
@@
To study the emergence of dynamical patterns, we introduce the local complex parameter
@@
When all particles are cycling in synchrony, the amplitude $r_{\bf j}\approx 1$ and the phase $\psi_{\bf j}$ are homogeneous. The period of $\psi_{\bf j}$ increases with $\varepsilon$, and eventually diverges. Before diverging, it undergoes large temporal fluctuations, which may desynchronise nearby sites, thus promoting spatial fluctuations of $\psi_{\bf j}$. At large $D_{0}$, spatial fluctuations are suppressed by the rapid displacement of particles in the system. Instead, at moderate $D_{0}$, such fluctuations can potentially build up into a large-scale instability.
Figure 2: (a) Snapshots of the local phase $\psi_{j}$ [Eq. (3), color code in Fig. 1(d)]. (b) Trajectories of the synchronisation parameter $R(t)$ [Eq. (4)]. (c) The averaged parameter $\langle\bar{R}\rangle$ and (d) its standard deviation $\sigma_{R}$ [Eq. (5)] lead to guidelines delineating boundaries between four phases. The solid and dashed lines correspond respectively to $\langle\bar{R}\rangle=0.2$ and $\sigma_{R}=0.045$. Symbols refer to the parameter values taken for panels (a-b) and Fig. 3. $(L,\rho_{0})=(128,10)$, and $f=\varepsilon/2$.
Instabilities can lead to the spatial coexistence of three cycling domains where $r_{\bf j}\equiv r_{{\bf j},c}\approx 1$. Interfaces between domains have a finite width where $r_{{\bf j},c}>r_{\bf j}>0$. The meeting points of interfaces are given by defects where $r_{\bf j}\approx 0$. Since all domains cycle at the same frequency, defects effectively rotate, thus forming spiral waves with a threefold symmetry [Figs. 1(c-d)]. Defects connected by the same interface rotate in opposite directions, and can annihilate by pair when colliding. Higher $D_{0}$ increases the domain sizes and interface widths, thus reducing the number of defects. Higher $\varepsilon$ reduces the interface widths and increases $r_{{\bf j},c}$ [Fig. 2(a)].
Remarkably, we do not observe any phase with turbulent waves, in contrast with the case of continuously deformable particles [27]. Such a turbulence stems in [27] from excitations of the homogeneous arrested phase which promote some localized, aperiodic cyclings. The discrete symmetry of PLM, which entails three arrested states, prevents such events by trapping the phase before it completes one cycle. Therefore, waves in PLM spontaneously organize into steady spirals. Note that the merging of defects can actually also stabilize planar waves. Besides, when initially ordered, the system can also accommodate circular waves without any defect [46].
Figure 3: Probability density of the synchronisation parameter $R$ [Eq. (4)] across the phase boundaries in Fig. 2. (a) The arrested-wave transition and (b) the cycling-wave transition both display a regime of metastability. Distribution measured over 1024 realizations.
@@
We examine how the distribution $P(R)$ varies across transitions [Fig. 3]. Deep in the wave and arrested phases [resp. pink and dark blue lines], $P(R)$ has a single peak, respectively at small and large $R$. Deep in the cycling phase [green line], $P(R)$ is nonzero for a finite domain of $R$ and peaks at its boundaries, due to the oscillations of $R(t)$. Going from waves to arrest, $P(R)$ becomes bimodal, thus signalling metastability [Fig. 3(a)]. Similarly, going from waves to cycling, $P(R)$ is now nonzero in two separated domains, due to the coexistence between the two dynamical states [Fig. 3(b)]. In short, both the arrest-wave and cycling-wave transitions feature a metastable regime.
Gauge invariance in the hydrodynamics.–Although spiral waves have been already reported in many RDS, one may wonder how the patterns in Figs. 1(c-d) actually differ from standard instabilities present, for instance, in the CGLE [30]. To address this question, we coarse-grain PLM in terms of $A_{\bf j}(t)\to{\cal A}({\bf x},t)$ in the continuum limit. Neglecting density fluctuations, and expanding to lowest orders in ${\cal A}$, we get
@@
${\cal L}({\cal A})$ $=c_{1}{\cal A}+c_{2}{{\cal A}^{*}}^{2}+c_{3}|{\cal A}|^{2}{\cal A},$
where $*{}^{*}$ refers to complex conjugation, $\Lambda$ is a zero-mean Gaussian white noise, $D\propto D_{0}$ is the macroscopic diffusion coefficient, and $(c_{1},c_{2},c_{3})$ are complex coefficients fixed by $(f,\varepsilon)$ [46]. The ${{\cal A}^{*}}^{2}$ term in Eq. (6) breaks the continuous gauge ${\cal A}\to{\cal A}\,e^{i\phi}$, in contrast with the standard CGLE [30]. Yet, this term preserves the discrete gauge ${\cal A}\to{\cal A}\,e^{\frac{2\pi i}{3}k}$, in contrast with the hydrodynamics of continuously deformable particles [27]. To the third order in ${\cal A}$, Eq. (6) actually contains all the terms compatible with this gauge.
@@
The local hydrodynamics $\partial_{t}{\cal A}={\cal L}({\cal A})$ reproduces the three homogeneous phases of PLM (disorder, cycles, and arrest) stemming from the interplay between the active drive and the energy landscape [Figs. 4(a-c)]. The breakdown of the continuous gauge ensures the existence of the arrested phase, as in [27]. The discrete gauge enforces that the local hydrodynamics can actually arrest in either one of three states, at variance with [27]. This degeneracy directly affects the shape of the hydrodynamic patterns. Indeed, Eq. (6) entails an instability promoting the spatial coexistence of cycling domains. Rotating defects with threefold symmetry spontaneously form where interfaces meet [Figs. 4(d-e)], yielding the same spiral waves as in PLM [Figs. 1(c-d)].
@@
In short, our coarse-graining shows that PLM are not captured by the standard CGLE [30], and therefore differ from standard RDS. It is striking that, although our hydrodynamics is an expansion at small $\cal A$, the lowest-order nonlinearities are here sufficient to reproduce the specific shape of particle-based spirals. Through the discrete gauge invariance, these nonlinearities bear the hallmark of the microscopic discrete symmetry.
Figure 4: Stream plot of the local hydrodynamics $\partial_{t}{\cal A}=\mathcal{L}({\cal A})$ [Eq. (6)] for (a) disorder, (b) cycles, and (c) arrest. Symbols indicate stable fixed points (green circles), unstable fixed points (red circles), and saddles (yellow squares). (d-e) Spiral waves centered around rotating defects spontaneously form in the hydrodynamics, analogously to the patterns in PLM [Figs. 1(c-d)]. $(L,f,\varepsilon,D,\rho_{0})=(512,0.5,2.5,100,10)$.
@@
Discussion.—We reveal that PLM entail spiral waves stemming from the competition between arrest and synchronisation. Therefore, our results show that the arrest-synchronisation scenario for pattern formation [27] extends to a broad class of models with discrete symmetry. Such a symmetry suffices to promote arrest, thus providing a mechanism distinct from that of repulsive deformable particles [27]. Waves are shaped by the discrete gauge invariance at the hydrodynamic level, which distinguishes them from the patterns of other RDS [31, 32, 33].
@@
Our work paves the way to examining the interplay between discrete symmetry and pattern formation for arbitrary $q$. Indeed, we expect that the three homogeneous phases (disorder, cycles, arrest) are robust beyond $q=3$. To study patterns, one can use a top-down approach postulating the hydrodynamics by identifying the terms which obey the discrete gauge ${\cal A}\to{\cal A}e^{\frac{2\pi i}{q}k}$. As in our coarse-graining [46], one can also derive the hydrodynamic coefficients in terms of the microscopic parameters. It is tempting to speculate that, depending on the parity of $q$, such a hydrodynamic study could lead to identifying generic properties of defect dynamics, and potential connections with the topology of other active models [47].

### v3

# Diffusive oscillators capture the pulsating states of deformable particles

@@ Abstract
We study a model of diffusive oscillators whose internal states are subject to a periodic drive. These models are inspired by the dynamics of deformable particles with pulsating sizes, where repulsion leads to arrest the internal pulsation at high density. We reveal that, despite the absence of any repulsion between the diffusive oscillators, our model still captures the emergence of dynamical arrest. We demonstrate that arrest here stems from the discrete nature of internal states which enforces an effective energy landscape analogous to that of deformable particles. Moreover, we show that the competition between arrest and synchronisation promotes spiral waves reminiscent of the pulsating states of deformable particles. Using analytical coarse-graining, we derive and compare the collective dynamics of diffusive oscillators with that of deformable particles. This comparison leads to rationalize the emergence of spirals in terms of a rotational invariance at the coarse-grained level, and to elucidate the role of hydrodynamic fluctuations.
@@
Introduction.—Active matter encompasses systems where a constant energy injection at the particle level leads to collective dynamics far from equilibrium, as is the case for many chemical, biological, and robotic systems [1, 2, 3, 4]. In the last decades, most studies of active matter have focused on the role of self-propulsion, i.e. the ability of each particle to independently undergo directed motion. This paradigm has led to the theoretical understanding of several kinds of collective dynamics which have no counterpart in equilibrium systems [5, 6].
@@
The energy injection in complex units need not result only in a self-propulsion mechanism. An important example is the case of deformable particles [7]. Indeed, complex aggregates such as macromolecules or living cells can change their shape due to internal activity [8], which leads to the spontaneous propagation of contraction waves in dense tissues. Such a wave propagation plays a crucial role in morphogenesis [9, 10, 11, 12, 13, 14], uterine contraction [15, 16], and cardiac arrhytmogenesis [17, 18].
@@
Vertex models are popular to capture the behavior of dense active systems such as biological tissues [19]. They typically consider self-propulsion as the only active component and investigate how it affects the rigidity transition [20, 21]. Yet, when dense tissues behave like solids, it is questionable whether self-propulsion should be the key ingredient. Other models have considered dense assemblies of deformable particles [22, 23, 24, 25, 26, 27, 28, 29, 30], where energy injection occurs through the sustained oscillation of individual sizes. With a Kuramoto-like synchronisation [31, 32] of particle sizes, contraction waves spontaneously emerge [26, 27, 33], which are reminiscent of those reported in biological systems [13, 17].
Contraction waves in deformable active particles stem from the competition between synchronisation and steric repulsion [27, 33]. The former favors a global cycling of particle sizes. The latter favors some specific deformations, which is analogous to enforcing an external potential through which particle sizes are driven [Fig. 1(a)]. At high density, local minima are deep enough to trap the particle size and thus promote an arrested state. Here, arrest refers to the hampering of size cycling, whereas it corresponds to the hampering of particle displacement within the rigidity transition of [20].
Figure 1: (a) The collective dynamics of deformable particles maps into a continuous oscillator $\psi$ evolving in a potential $U$ and subject to a drive $f$ [34, Sec. S1.D]. (b) The internal dynamics of our discrete oscillator follows some transitions biased by the drive $f$ without any potential. (c) Our oscillators freely diffuse at a rate $D_{0}$ without any volume exclusion.
@@
Interestingly, there are alternative mechanisms for arrest and synchronisation beyond steric repulsion and Kuramoto-like interactions. This is a motivation for exploring whether these alternatives lead to a phenomenology similar to that of deformable active particles. Indeed, one may wonder whether the arrest-synchronisation competition is actually a generic scenario for pattern formation. If so, it encourages one to search for any hydrodynamic invariance and/or broken symmetry which might stand out as a hallmark of this competition.
@@
In this Letter, we study a diffusive oscillator model (DOM) [Fig. 1] featuring spiral waves rotating around defects [Fig. 2]. We reveal that the discreteness of internal states is key to arresting the dynamics, which, in combination with synchronisation and drive, generically yields wave propagation. Internal states can here be regarded as representing the local mimina of an external potential [Fig. 1(a)], by analogy with the case of deformable particles [27, 33]. Through analytical coarse-graining, we map our DOM into a specific form of the complex Ginzburg-Landau equation (CGLE) [35] which breaks the continuous rotational invariance, while maintaining a discrete rotational invariance. We argue that such an invariance not only signals the emergence of arrest, but also constrains the features of the waves at the hydrodynamic and microscopic levels. Moreover, we reveal that hydrodynamic noise is essential to forming patterns, whereas density fluctuations are irrelevant.
Overall, our results demonstrate that the arrest-synchronisation competition, present in our DOM and in deformable particles [27, 33], is a generic route for forming patterns which are distinct from those of standard reaction-diffusion systems (RDS) [36, 37, 38].
@@
Diffusive oscillator model.—We consider $N$ oscillators in $V=L^{d}$ sites of a hypercubic $d$-dimensional lattice, with global number density $\rho_{0}=N/V$, without any excluded volume. Each site contains an arbitrarily large number of oscillators. Each oscillator has an internal state, labeled by a discrete index $a\in\{1,\ldots,q\}$, as a proxy to mimic the internal phase of deformable particles [22, 23, 24, 25, 26, 27, 33]. The crucial difference is that such states now feature a discrete symmetry. In what follows, we focus on the case $q=3$, which is the minimum number of states to accommodate a current, and $d=2$. The system configuration $\{n_{{\bf j},a}\}$ is then given by the number of oscillators for each state $a$ and site ${\bf j}$.
@@
At every time step $dt/N$, an oscillator with state $a$ can either jump to a neighbouring site with probability $D_{0}\,dt$, or switch to state $b$ with probability $W_{ba}\,dt$ [Figs. 1(b,c)]:
@@
The first term in the exponent of Eq. (1) is defined by $f_{ab}=\pm f$ when $a-b=\pm 1\mod 3$, and $f_{ab}=0$ otherwise. This drive mimics the pulsation of deformable particles [22, 23, 24, 25, 26, 27, 33]. Inspired by some active lattice dynamics [39, 40, 41, 42, 43, 44, 45], the second term accounts for a synchronising Potts-like [46] interaction with strength $\varepsilon$, where $\rho_{\bf j}=\sum_{a}n_{{\bf j},a}$ is the local density at site $\bf j$. This term favors transition towards the state with the highest number of oscillators locally.
@@
While interactions between oscillators are fully connected onsite, different sites exchange information only via diffusion. At small $D_{0}$, many transitions of internal states occur in between two rare jumps of sites. At large $D_{0}$, jumps are so frequent that now all oscillators are effectively interacting in between two state transitions. In both cases, patterns cannot emerge, and the system can be described solely in terms of the onsite dynamics [47, 48].
Figure 2: The amplitude $r$ and the phase $\psi$ of synchronization [Eq. (5)] reveal the coexistence of ordered domains in (a-b) microscopic dynamics [Eq. (1)], and (c-d) hydrodynamics [Eq. (8)]. Rotating defects appear at the meeting point of domain interfaces, forming spiral waves with a threefold symmetry. (a) $(L,f,\varepsilon,D_{0},\rho_{0})=(512,1.25,2.5,10^{3},10)$, and (b) $(L,f,\varepsilon,D,\rho_{0})=(512,0.5,2.5,100,10)$.
Breakdown of rotational invariance: Analogy with deformable particles.—In the absence of diffusion ($D_{0}=0$), the evolution of the onsite occupation numbers $n_{a}$ can be studied in terms of the collective complex variable
$A(t)=\frac{1}{\rho_{0}}\sum_{a=1}^{3}e^{\frac{2\pi i}{3}a}n_{a}(t)\equiv r(t)\,e^{i\psi(t)}.$ (2)
Coarse-graining the microscopic dynamics, and expanding to the lowest orders in $A$, we get [34, Secs. S1.A-B]
$\dot{A}=c_{1}A+c_{2}{A^{*}}^{2}+c_{3}|A|^{2}A\equiv{\cal L}(A),$ (3)
where refers to complex conjugation, and $(c_{1},c_{2},c_{3})$ are complex coefficients fixed by $(f,\varepsilon)$ [34, Sec. S1.B]. In the regime of large $\rho_{0}$, fluctuations become irrelevant. The ${A^{*}}^{2}$ term in Eq. (3) breaks the continuous symmetry $A\to A\,e^{i\phi}$, in contrast with the standard Stuart-Landau oscillator [49]. Yet, this term preserves the discrete symmetry $A\to A\,e^{\frac{2\pi i}{3}k}$ ($k$ integer). To the third order in $A$, Eq. (3) actually contains all the terms compatible with this rotational symmetry.
@@
From the dynamics in Eq. (3), it follows that there are three stable phases [47, 48][Fig. 3]: (i) a disordered phase at small $\varepsilon$, where oscillators are uniformly distributed in the three states, with symmetric fixed point $n_{a}=\rho_{0}/3$, and $|A|$ vanishes at large $\rho_{0}$; (ii) a cycling phase at intermediate $\varepsilon$, where oscillators collectively undergo periodic transitions between states; (iii) an arrested phase at large $\varepsilon$, with three fixed points invariant under the cyclic permutations $A\to A\,e^{\frac{2\pi i}{3}k}$.
Figure 3: Stream plots of the collective variable $A$ [Eq. (3)] for (a) disorder, (b) cycles, and (c) arrest, with stable fixed points (green circles), unstable fixed points (red circles), and saddles (yellow squares). (d) Effective landscape $U_{\rm eff}$ of the collective phase $\psi$ [Eq. (4)] with $f=0.1$. (e) Stochastic trajectories of $\psi$ measured in the onsite dynamics [Eq. (1)]. As $\varepsilon$ increases, phase oscillations slow down, and eventually reach arrest. $(N,f)=(10^{3},2)$ and $\varepsilon$ as in (d). The natural pulsation reads $\omega=\sqrt{3}\sinh f$ [34, Sec. S1.B].
The breakdown of the continuous symmetry $A\to A\,e^{i\phi}$ ensures the existence of arrest, while the discrete symmetry $A\to A\,e^{\frac{2\pi i}{3}k}$ enforces that the arrested phase is actually degenerate. Note that, in the microscopic model [Eq. (1)], transitions between states can be regarded as unimolecular reactions, in contrast with some models of multimolecular RDS [50, 51], and the total number of oscillators is conserved. These features lead to stabilising disorder at small $\varepsilon$, and ensure that arrest at large $\varepsilon$ is not an absorbing phase.
In the arrested phase, the discrete nature of the internal states enforces that the collective phase $\psi$ is subject to the effective landscape $U_{\rm eff}$ [34, Sec. S1.C]:
$\dot{\psi}=-dU_{\rm eff}/d\psi.$ (4)
Remarkably, $U_{\rm eff}$ features a series of local periodic minima whose depth increases with $\varepsilon$ [Fig. 3(d)], which can be rationalized in simple terms. At large $\varepsilon$, while continuous oscillators [31] synchronously cycle without any cost, the collective cycling of our discrete oscillators [Eq. (1)] entails some periodic desynchronisations: the minima in $U_{\rm eff}$ describe the cost of such desynchronisations. As $\varepsilon$ increases, transitions between minima are less favored, so that our oscillators spend more time in a given state before cycling to the next one. Above a critical $\varepsilon$, this effect completely counteracts the drive, breaking down the periodicity of oscillations, and eventually stabilizing arrest [Fig. 3(e)].
Interestingly, a similar phase trapping has been reported in deformable particles [26, 27, 33]. Here, repulsion between particles is equivalent to an external potential constraining the deformation statistics [Fig. 1(a)]. At the collective level, such a constraint can be recapitulated in terms of a landscape with periodic minima [34, Sec. S1.D], qualitatively analogous to $U_{\rm eff}$ [Fig. 3(d)]. In that respect, our DOM captures the same phenomenology, with identical collective states (disorder, arrest, cycles) as that of pulsating deformable particles.
@@
In short, arrest emerges in our DOM solely due to the discreteness of internal states. As a result, our DOM entails a competition between arrest and synchronisation despite the absence of any repulsion, in contrast with deformable particles [26, 27, 33]. This competition opens the door to the emergence of dynamical patterns in spatially extended systems.
@@
Spiral waves rotate around defects.—In the presence of diffusion ($D_{0}>0$), the displacement of oscillators follows a free dynamics, independently of any interaction, so that the density profile is always homogeneous. Yet, the spatial distribution of the oscillator states may not remain homogeneous. Indeed, even when oscillators are synchronised on site, they might not be synchronised between sites, so that the system can potentially accommodate spatial instabilities.
@@
To study the emergence of dynamical patterns, we introduce the local complex variable
@@
When all oscillators are cycling in synchrony, the amplitude $r_{\bf j}\approx 1$ and the phase $\psi_{\bf j}$ are homogeneous. The period of $\psi_{\bf j}$ increases with $\varepsilon$, and eventually diverges. Before diverging, it undergoes large temporal fluctuations, which may desynchronise nearby sites, thus promoting spatial fluctuations of $\psi_{\bf j}$. At large $D_{0}$, spatial fluctuations are suppressed by the rapid displacement of oscillators in the system. Instead, at moderate $D_{0}$, such fluctuations can potentially build up into a large-scale instability.
Figure 4: (a) Snapshots of the local phase $\psi_{j}$ [Eq. (5), color code in Fig. 2]. (b) Trajectories of the synchronisation parameter $R(t)$ [Eq. (6)]. (c) The averaged parameter $\langle\bar{R}\rangle$ and (d) its standard deviation $\sigma_{R}$ [Eq. (7)] lead to guidelines delineating boundaries between four phases. The solid and dashed lines correspond respectively to $\langle\bar{R}\rangle=0.2$ and $\sigma_{R}=0.045$. Symbols refer to the parameter values taken for panels (a-b) and (e-d). (e,f) Distribution of $R$ for various phases (1024 realizations). The arrested-wave and cycling-wave transitions both display metastability. $(L,\rho_{0})=(128,10)$, and $f=\varepsilon/2$.
Instabilities can lead to the spatial coexistence of three cycling domains where $r_{\bf j}\equiv r_{{\bf j},c}\approx 1$. Interfaces between domains have a finite width where $r_{{\bf j},c}>r_{\bf j}>0$. The meeting points of interfaces are given by defects where $r_{\bf j}\approx 0$. Since all domains cycle at the same frequency, defects effectively rotate, thus forming spiral waves with a threefold symmetry [Figs. 2(a-b)]. Defects connected by the same interface rotate in opposite directions, and can annihilate by pairs when colliding. Higher $D_{0}$ increases the domain sizes and interface widths, thus reducing the number of defects. Higher $\varepsilon$ reduces the interface widths and increases $r_{{\bf j},c}$ [Fig. 4(a)].
Remarkably, we do not observe any states with turbulent waves, in contrast with the case of continuously deformable particles [27]. Such a turbulence stems in [27] from excitations of the homogeneous arrested phase which promote some localized, aperiodic cyclings. The discrete symmetry of our DOM, which entails three arrested states, prevents such events by trapping the phase before it completes one cycle. Therefore, waves in our DOM spontaneously organize into steady spirals with threefold symmetry. Note that the merging of defects can actually also stabilize planar waves. Besides, when initially ordered, the system can also accommodate circular waves without any defect [34, Sec. S2.A].
@@
We examine how the distribution $P(R)$ varies across transitions. Deep in the wave and arrested phases [resp. pink and dark blue lines], $P(R)$ has a single peak, respectively at small and large $R$. Deep in the cycling phase [green line], $P(R)$ is nonzero for a finite domain of $R$ and peaks at its boundaries, due to the oscillations of $R(t)$. Going from waves to arrest, $P(R)$ becomes bimodal, thus signaling metastability [Fig. 4(e)]. Similarly, going from waves to cycling, $P(R)$ is now nonzero in two separate domains, due to the coexistence between the two dynamical states [Fig. 4(f)]. In short, both the arrest-wave and cycling-wave transitions feature a metastable regime.
The role of hydrodynamic fluctuations.–Although spiral waves have already been reported in many RDS, one may wonder how our patterns [Figs. 2(a-b)] actually differ from standard instabilities present, for instance, in the CGLE [35]. To address this question, we coarse-grain our DOM in terms of $A_{\bf j}(t)\to{\cal A}({\bf x},t)$ and $\rho_{\bf j}(t)\to\rho({\bf x},t)$ in the continuum limit, and expand to the lowest orders in ${\cal A}$ [34, Sec. S1.B]:
@@
where $\cal L$ is defined in Eq. (3), and $D\propto D_{0}$ is the macroscopic diffusion coefficient. Neglecting the fluctuations of the coarse-grained density $\rho({\bf x},t)$, we deduce that it obeys the simple diffusion equation ($\partial_{t}\rho=D\nabla^{2}\rho$) independently of $\cal{A}$, and relaxes to the homogeneous profile $\rho=\rho_{0}$. Thus, we approximate $\Lambda$ as an additive zero-mean Gaussian white noise with correlations proportional to $1/\rho_{0}$ [34, Sec. S1.B].
@@
The degeneracy of the arrested state directly affects the shape of the hydrodynamic patterns. Indeed, Eq. (8) entails an instability promoting the spatial coexistence of cycling domains. Rotating defects with threefold symmetry spontaneously form where interfaces meet [Figs. 2(c-d)], yielding the same spiral waves as in our DOM [Figs. 2(a-b)]. Remarkably, in the absence of noise ($\Lambda=0$), the homogeneous states are always stable [34, Sec. S1.B]. In other words, while density fluctuations can be safely neglected, the fluctuations in the hydrodynamics of $\cal A$ are essential to capture patterns, as in [27].
@@
In short, our coarse-graining shows that the hydrodynamics of our DOM is distinct from the standard CGLE [35], so that our DOM clearly differs from standard RDS. In practice, adding to the standard CGLE the lowest-order nonlinearities, compatible with the discrete symmetry, suffices to reproduce the specific shape of spirals observed in our DOM.
@@
Discussion.—Our DOM with discrete symmetry entails spiral waves stemming from the competition between arrest and synchronisation [Fig. 2]. The key idea is that discrete states enforce an effective landscape [Fig. 3(d)] equivalent to the case of deformable particles with repulsive interactions [34, Sec. S1.D][27, 33]. Discreteness of states here suffices to promote arrest, thus providing a mechanism distinct from that at play in [27, 33]. Therefore, our results show that the arrest-synchronisation scenario for pattern formation extends to a broad class of models with discrete symmetry. Remarkably, waves are shaped by the discrete rotational invariance at the hydrodynamic level, which distinguishes them from the patterns of other RDS [36, 37, 38]. Besides, the fluctuations of the complex field are crucial to yield patterns, whereas density fluctuations are irrelevant.
@@
Our work paves the way to examining the interplay between discrete symmetry and pattern formation for arbitrary $q$. Indeed, we expect that the three homogeneous phases (disorder, cycles, arrest) are robust beyond $q=3$. To study patterns, one can use a top-down approach postulating the hydrodynamics by identifying the terms which obey the discrete symmetry ${\cal A}\to{\cal A}e^{\frac{2\pi i}{q}k}$. As in our coarse-graining [34, Sec. S1], one can also derive the hydrodynamic coefficients in terms of the microscopic parameters. It is tempting to speculate that, depending on the parity of $q$, such a hydrodynamic study could lead to identifying generic properties of defect dynamics [52, 53], and potential connections with the topology of other active models with phase trapping [45, 54, 55, 56, 57].